OpenStudy (ammarah):

What term is needed to add to each side to complete the square? x2x^+5x=7

OpenStudy (ammarah):

@ganeshie8

OpenStudy (ammarah):

i divided it by 2 all the coefficents

OpenStudy (ammarah):

sorry theres no x in the front

OpenStudy (ammarah):

just 2x^2

ganeshie8 (ganeshie8):

\(\large 2x^2+5x = 7\)

ganeshie8 (ganeshie8):

like this ?

OpenStudy (ammarah):

yes

ganeshie8 (ganeshie8):

yes, after dividing, just take half of the x coefficient, and square it

ganeshie8 (ganeshie8):

\(\large x^2+\frac{5}{2}x = \frac{7}{2}\)

ganeshie8 (ganeshie8):

whats half of \(\large \frac{5}{2}\) ?

OpenStudy (ammarah):

2.5

ganeshie8 (ganeshie8):

half of \(\large \frac{5}{2}\) is \(\large \frac{5}{2}/2 = \frac{5}{4}\)

ganeshie8 (ganeshie8):

square it and add both sides

OpenStudy (ammarah):

oh lol

OpenStudy (ammarah):

but squaring it gives a decimal

ganeshie8 (ganeshie8):

:) \(\large x^2+\frac{5}{2}x + \left(\frac{5}{4}\right)^2 = \frac{7}{2} + \left(\frac{5}{4}\right)^2\)

ganeshie8 (ganeshie8):

left hand side can be written as a perfect square : a^2 + 2ab + b^2 = (a+b)^2

ganeshie8 (ganeshie8):

and, \(\large \left(\frac{5}{4}\right)^2 \) is the term we needed to add both sides in order to do that

ganeshie8 (ganeshie8):

see if that looks okay...

OpenStudy (ammarah):

ok

OpenStudy (ammarah):

what then?

ganeshie8 (ganeshie8):

we're done

ganeshie8 (ganeshie8):

simplify

ganeshie8 (ganeshie8):

\(\large \left(\frac{5}{4}\right)^2\) \(\large \frac{25}{16} \) is the term to be added both sides,

ganeshie8 (ganeshie8):

oh wait a sec, u need to multiply 2 both sides cuz u have only 7 on rihgt hand side initially

ganeshie8 (ganeshie8):

so the term to be added is : \(\large \frac{25}{8} \) is the term to be added both sides

ganeshie8 (ganeshie8):

u have options for this ?

OpenStudy (ammarah):

ok how about 2x^2-32x=56 i solved it and got x^-16=28

ganeshie8 (ganeshie8):

\(\large 2x^2 -32x = 56\) divide by 2 both sides : \(\large x^2 -16x = 28\)

ganeshie8 (ganeshie8):

take half of "x" coefficient, and add its square both sides

ganeshie8 (ganeshie8):

half of 16 is 8, so add 8^2 both sides

ganeshie8 (ganeshie8):

\(\large 2x^2 -32x = 56\) divide by 2 both sides : \(\large x^2 -16x = 28\) add 8^2 both sides : \(\large x^2 -16x +8^2= 28 + 8^2\)

OpenStudy (ammarah):

ok is that my final equation?

OpenStudy (ammarah):

so 64

ganeshie8 (ganeshie8):

yes, left side can be written as perfect square. so we're done

ganeshie8 (ganeshie8):

yes, its either 64 or 128

ganeshie8 (ganeshie8):

depends on whatever options you're given

ganeshie8 (ganeshie8):

cuz, earliear u have divided 2 both sides right ?

ganeshie8 (ganeshie8):

so to account for that u need to multiply 2 both sides again

ganeshie8 (ganeshie8):

so, answer can be either 64 or 128

ganeshie8 (ganeshie8):

whatever exists in ur options, tick that

OpenStudy (ammarah):

ok....